OpenStudy (lilai3):

PLEASE HELP!!! WILL FAN/REWARD MEDAL/ WRITE TESTIMONY, ETC. ETC.!! A river is rising at a rate of 3 inches per hour. If the river rises more than 2 feet, it will exceed flood stage. How long can the river rise at this rate without exceeding flood stage? Please explain. Thank you!~

OpenStudy (the_fizicx99):

3in/Hr > or equal to 24in/xHRs 24 inches in 2 feet

OpenStudy (anonymous):

2 foot =24 inches each hour river rises 3 inches time taken to reach 24 inches=24/3=8hours Did you get it??

OpenStudy (anonymous):

3in/Hr > or equal to 24in/xHRs 24 inches in 2 feet

OpenStudy (the_fizicx99):

That'ss your inequality. So overall l'd say 8 hours

OpenStudy (the_fizicx99):

Thank you @Jacob293 for copying my answer.

OpenStudy (anonymous):

;) <3 @tHe_FiZiCx99

OpenStudy (the_fizicx99):

No. @Jacob293

OpenStudy (anonymous):

I love you

OpenStudy (lilai3):

lol you guys are funny.

OpenStudy (lilai3):

ALL of you were great help! salmon explained it well.

OpenStudy (the_fizicx99):

@lilai3 replace x with 3.. so 24/3 = 8 So, 3in/Hr > or equal to 24in/3 = 8 hours

OpenStudy (lilai3):

yes, i know. i read salmon's explanation. thank you!

OpenStudy (anonymous):

No problem anytime :) xD @lilai3

OpenStudy (the_fizicx99):

^ Just smh

OpenStudy (anonymous):

SAME^